Sandy’s first four test scores this term are 73, 87, 81, and 76. To have at least a B
test grade, Sandy needs to average at least 80 on the five term tests (which count
equally). Let t represent Sandy’s score on the fifth test, and write an inequality that
describes the range of t-values that will meet Sandy’s goal.

Answer by Theo(13342)

the formula you need is:

(73 + 87 + 81 + 76 + t) / 5 >= 80

multiply both sides of this equation by 5 to get:

(73 + 87 + 81 + 76 + t >= 400

simplify to get:

317 + t >= 400

subtract 317 from both sides of this equation to get:

t >= 83

she has to get at least an 83 on the last test.

the range of values she can get are 83 <= t <= 100

100 is the upper limit because it is impossible to get more than that on any test unless they give extra credit.

extra credit to make a test score greater than 100 is not assumed.

the general equation for graphing is:

y = (317 + x) / 5

graph this equation and place a vertical line at x = 83 and x = 100 and the value of y between those vertical lines is the range of possible averages you will be able to get.

the graph is shown below:
your average will be a minimum of 80 and a maximum of 83.4.
the range of t values is betwee 83 and 100 inclusive.

$$$